OpenStudy (anonymous):

Mia walks with a velocity of 1.3 m/s south. She does so while riding on a train that is traveling with a velocity of 4.6 m/s north. What is Mia's velocity?

OpenStudy (anonymous):

A. 3.3 m/s north B. 1.3 m/s south C. 1.3 m/s north D. 3.3 m/s south

OpenStudy (raffle_snaffle):

I think it's D, but not 100% sure.

OpenStudy (anonymous):

4.6 north - 1.3 south = answer

OpenStudy (anonymous):

If you think of setting North as positive, and South as negative, then your velocities are \[v_1 = 4.6m/s \ N = +4.6m/s \\ v_2 = 1.3m/s \ S = -1.3m/s\] And the net velocity would be \[v_{net} = v_1+v_2\] if the result is positive then she's walking North, and if it's negative she's walking South.